# Programming with Pixels: Computer-Use Meets Software Engineering

### v1

@@ Abstract
Recent advancements in software engineering (SWE) agents have largely followed a tool-based paradigm, where agents interact with hand-engineered tool APIs to perform specific tasks. While effective for specialized tasks, these methods fundamentally lack generalization, as they require predefined tools for each task and do not scale across programming languages and domains. We introduce Programming with Pixels (PwP), an agent environment that unifies software development tasks by enabling computer-use agents—agents that operate directly within an IDE through visual perception, typing, and clicking, rather than relying on predefined tool APIs. To systematically evaluate these agents, we propose PwP-Bench, a benchmark that unifies existing SWE benchmarks spanning tasks across multiple programming languages, modalities, and domains under a task-agnostic state and action space. Our experiments demonstrate that general-purpose computer-use agents can approach or even surpass specialized tool-based agents on a variety of SWE tasks without the need for hand-engineered tools. However, our analysis shows that current models suffer from limited visual grounding and fail to exploit many IDE tools that could simplify their tasks. When agents can directly access IDE tools, without visual interaction, they show significant performance improvements, highlighting the untapped potential of leveraging built-in IDE capabilities. Our results establish PwP as a scalable testbed for building and evaluating the next wave of software engineering agents. [^1] Carnegie Mellon University. We release code and data at programmingwithpixels.com.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Comparison between the traditional tool-based paradigm (left) and our proposed Programming with Pixels (PwP) framework (right) for software engineering (SWE) agents. Instead of relying on specialized hand-engineered tools, PwP enables agents to interact directly with an IDE through basic computer interactions and screen observations. The framework naturally allows agents to take advantage of existing IDE capabilities across multiple languages and tools. Furthermore, PwP-Bench is a comprehensive benchmark to evaluate agent performance across different SWE domains.
Human software developers possess a remarkable ability to work across a wide range of programming tasks, seamlessly adapting to new languages, tools, and problem domains. Realizing a single, general-purpose agent with similar versatility is the overarching goal of many recent efforts in code generation and software engineering automation (Jiang et al., 2024; Jin et al., 2024; Wang et al., 2024b). However, most software engineering agents (SWE agents) still rely on a tool-based paradigm, where an agent takes actions using hand-engineered functions (e.g., search repository, run Python code) exposed through a text API (Yang et al., 2024a, b; Wang et al., 2024a, b). This fundamentally limits generalization, since tool-based agents can only perform tasks using the predefined actions. For example, an agent designed to manage GitHub pull requests lacks debugging abilities unless it is programmed into the agent’s API. Furthermore, the tool-based paradigm lacks scalability, as hand-engineering complex tools requires significant human effort and may not be bug-free. As a result, it remains unclear whether the tool-based paradigm scales to the diversity of software engineering tasks, which spans multiple languages, modalities, and task types.
Our motivating hypothesis is that achieving general-purpose SWE agents requires a shift to computer-use agents (Anthropic, 2024) that interact with computers as humans do: by observing the screen, typing, and clicking. To this end, we recast agentic software engineering as interacting directly with an integrated development environment (IDE) by observing its visual state and using basic actions such as clicking and typing. This allows the agent to perform any task possible in an IDE and leverage all of the IDE’s tools—from debuggers to web browsers—without requiring specialized APIs. However, despite promising results in web navigation (Anthropic, 2024) and open-ended computer tasks (Xie et al., 2024), the ability of computer-use agents to perform software engineering remains underexplored and we lack a dedicated environment for software engineering.
To close this gap, we introduce Programming with Pixels (PwP), the first software engineering agent environment aimed at general-purpose computer-use agents. The PwP environment is a VSCode-based IDE where agents perceive the screen and use primitive actions such as typing, pointing, and clicking. PwP fulfills two key properties. First, the environment is expressive, allowing agents to complete any software engineering task achievable in an IDE, without language- or domain-specific modifications. Second, agents naturally interact with any tools available in the IDE–including debuggers, linters, and code suggestions while handling diverse data types such as images, videos, and PDFs–through basic actions such as clicking and typing. This notion of tool use fundamentally differs from hand-engineered tool APIs, offering scalability and reducing the effort needed to hand-engineer tools for AI agents. Namely, PwP lets agents take advantage of the rich tools already accessible to humans in an IDE, rather than reinventing the wheel. Finally, computer-use agents reduce the need for complex tool pipelines (e.g., tool-specific prompts), opening up a simplified approach to general-purpose SWE agents.
To further evaluate agents developed for computer-use we construct PwP-Bench, a unified benchmark of 15 tasks spanning a variety of software engineering activities, including code generation, pull request resolution, UI development, and DevOps workflows. We show for the first time that general-purpose computer-use agents achieve non-trivial performance on a wide variety of SWE tasks, often approaching or exceeding state-of-the-art tool-based agents.
However, our analysis reveals substantial opportunities for future work. First, even state-of-the-art computer-use agents suffer from visual grounding issues. Second, we show that current agents lack the ability to use many of the tools available in the IDE, including ones that could make their tasks trivial. This suggests that training agents to explore and use the tools available in the IDE is a fruitful future direction. Finally, we find that only one model, Claude (Anthropic, 2024), performs well, highlighting the need for further research into training and improving computer-use agents.
@@
In summary, our contributions are as follows. First, we introduce Programming with Pixels (PwP), the first software engineering-focused environment for evaluating computer-use agents. Second, we introduce PwP-Bench, a benchmark spanning 15 diverse SWE domains, allowing for systematic comparison of software engineering agents. Third, we demonstrate, for the first time, that computer-use agents can perform a wide variety of software engineering tasks without additional hand-engineering of their action or observation spaces. Fourth, we analyze the limitations of current computer-use agents, identifying the need for models that better leverage IDE tooling and highlighting agent training as a key future direction. Finally, both existing agents and benchmarks can be easily incorporated into our unified environment, positioning PwP to serve as a common platform for developing future SWE agents. Overall, PwP challenges the prevailing tool-based paradigm for SWE agents and provides a platform for developing more general agents that interact directly with IDEs.
@@ 2 Related Work
2.1 Task-specific SWE benchmarks
Early neural code generation approaches were typically evaluated on fixed input-output pairs—for example, generating code from docstrings (Chen et al., 2021) or from general textual descriptions (Austin et al., 2021). Subsequent benchmarks extended these evaluations to interactive settings, such as resolving GitHub pull requests or writing unit tests for real-world code repositories (Jimenez et al., 2023; Zan et al., 2024; Mündler et al., 2025). More recently, efforts have broadened the scope of code generation to include multimodal tasks, where vision models must interpret images to generate correct code or edits (Si et al., 2024; Shi et al., 2024; Jing et al., 2024; Yang et al., 2024b). However, each of these benchmarks is confined to specific languages, modalities, or task types. In contrast, our proposed PwP-Bench unifies these diverse evaluations into a single framework, encompassing multimodal and multilingual challenges that require interaction with a broad suite of IDE tools. Using this unified approach we reproduce the performance of established benchmarks and encourage the development of general-purpose agents capable of handling a variety of new software engineering tasks. We further compare our work with previous efforts in Tables 1 and 2.
2.2 Software Engineering (SWE) Agents
2.3 Visual Agents and Computer-Use Agents
A family of recent multimodal agent benchmarks require agents to operate user interfaces using a predefined, limited set of actions (e.g., new_tab, go_back, click [element id](Koh et al., 2024a; Deng et al., 2023; Zheng et al., 2024) . These visual agents typically rely on additional prompting—such as set-of-marks techniques that supply an HTML accessibility tree containing textual and positional information—to overcome their inherent poor visual grounding capabilities (Yang et al., 2023b). Despite such aids, these agents often fail when faced with the complex and dense IDE interfaces found in our environment.
2.4 Expressive Agent Environments
Prior work on expressive agent environments has predominantly targeted the web domain (Koh et al., 2024a; Deng et al., 2023), entire operating systems (Xie et al., 2024; Bonatti et al., 2024; Rawles et al., 2023), or other general scenarios (Xu et al., 2024). Some of these environments, such as OSWorld (Xie et al., 2024), feature general action and observation spaces similar to ours. However, although these benchmarks are capable of expressing a wide range of tasks, they do not focus on the unique challenges inherent to software engineering within an IDE. For example, while OSWorld offers a broad set of tasks, it is not specifically designed for SWE, resulting in increased computational overhead. Software engineering is a diverse and important domain that merits its own dedicated environment.

### v2

# Programming with Pixels: Can Computer-Use Agents do Software Engineering?

@@ Abstract
Computer-use agents (CUAs) hold the promise of performing a wide variety of general tasks, but current evaluations have primarily focused on simple scenarios. It therefore remains unclear whether such generalist agents can automate more sophisticated and specialized work such as software engineering (SWE). To investigate this, we introduce Programming with Pixels (PwP), the first comprehensive computer-use environment for software engineering, where agents visually control an IDE to perform diverse software engineering tasks. To enable holistic evaluation, we also introduce PwP-Bench, a benchmark of 15 existing and new software-engineering tasks spanning multiple modalities, programming languages, and skillsets. We perform an extensive evaluation of state-of-the-art open-weight and closed-weight CUAs and find that when interacting purely visually, they perform significantly worse than specialized coding agents. However, when the same CUAs are given direct access to just two APIs—file editing and bash operations—performance jumps, often reaching the levels of specialized agents despite having a task-agnostic design. Furthermore, when given access to additional IDE tools via text APIs, all models show further gains. Our analysis shows that current CUAs fall short mainly due to limited visual grounding and the inability to take full advantage of the rich environment, leaving clear room for future improvements. PwP establishes software engineering as a natural domain for benchmarking whether generalist computer-use agents can reach specialist-level performance on sophisticated tasks.  [^1] Code and data released at https://programmingwithpixels.com
@@ 1 Introduction
Computer-use agents (CUAs) hold the promise of automating a wide range of economically valuable tasks by acting through primitive actions such as clicking, typing, and observing digital screens, potentially obviating the need for specialized AI agent action interfaces (Anthropic, 2024; OpenAI, 2025; Yang et al., 2024a). However, current evaluations have primarily focused on simple tasks such as web navigation (Koh et al., 2024), basic document editing, or tweaking settings in operating systems (Xie et al., 2024; Bonatti et al., 2024). Therefore, it remains unclear whether current generalist computer-use agents can automate more sophisticated and specialized tasks such as software engineering. In this work, we specifically study how well the current generation of computer-use agents can do software engineering and identify their key limitations.
The choice of using software engineering as the test domain is motivated by two primary reasons. First, software engineering represents an economically important and practically challenging task. Second, the field of AI software-engineering agents (SWE agents) has produced numerous specialized agents that use hand-engineered APIs for specific operations (Yang et al., 2024a; Wang et al., 2024b; Xia et al., 2024), providing strong baselines for comparison. These agents use custom functions such as file editing, code search, and repository management, with each tool requiring significant engineering effort and domain expertise. For instance, SWE-agent (Yang et al., 2024a) uses language-specific parsers and editing commands, while Agentless (Xia et al., 2024) relies on Python-specific abstract syntax trees. This specialization has yielded strong performance, but it raises a fundamental question: can general-purpose computer-use agents match specialized agents in complex domains like software engineering?
Figure 1: Programming with Pixels is an environment for computer-use agents, where they interact with a VSCode IDE through keyboard and mouse actions while observing the screen. The framework supports multiple programming languages, tests interactions with multiple IDE features, modalities (eg: text, images, data files). PwP-Bench evaluates agents across 15 diverse software engineering tasks such as code generation, UI generation, Data Science.
To investigate this question, we introduce Programming with Pixels (PwP), the first environment for systematically evaluating computer-use agents on software engineering tasks. The PwP environment provides a VSCode-based IDE where agents perceive the screen and use primitive actions such as typing and clicking to perform a variety of SWE tasks. This design enables two critical properties for fair evaluation. First, the environment is expressive, allowing agents to complete any software engineering task achievable in an IDE without language- or domain-specific modifications. Second, agents can access all IDE tools—debuggers, linters, code suggestions—through the same visual interface available to human developers or specialized SWE agents. Hence, PwP provides a general-purpose, realistic software engineering environment for testing computer-use agents.
To evaluate computer-use agents, we construct PwP-Bench, a benchmark of 15 tasks spanning different tasks such as code generation, pull request resolution, UI development, and data science across multiple programming languages and modalities. The benchmark represents a unification of 13 existing SWE tasks ported for evaluating computer-use agents, and 2 additional tasks developed by us. Our evaluation of state-of-the-art computer-use agents reveals that when restricted to pure visual interaction, these agents achieve only 22.9% average accuracy, significantly underperforming specialized coding agents. However, when augmented with just two basic text APIs—file editing and bash operations—the same agents achieve 50.7% accuracy, often approaching specialized agent performance despite their task-agnostic design. Furthermore, our analysis reveals substantial opportunities for future work. First, even state-of-the-art computer-use agents suffer from visual grounding issues. Second, we show that current computer-use agents lack the ability to use many of the tools available in the IDE, including ones that could make their tasks trivial. This suggests that training computer-use agents to explore and leverage the functionality present in their computer environment is a fruitful future direction. Overall, our results highlight software engineering as a realistic and challenging benchmark for evaluating and improving computer-use agents.
@@
In summary, our contributions are as follows. First, we introduce Programming with Pixels (PwP), the first software engineering-focused environment for evaluating computer-use agents. Second, we introduce PwP-Bench, a benchmark spanning 15 diverse SWE domains, allowing for systematic comparison of computer-use agents. Third, through extensive evaluation, we highlight the limitations of current computer-use agents, identifying the need for models that have better visual grounding and that better take advantage of their environment as key future directions. Finally, we open-source our environment and benchmark, allowing it to serve as an open platform for evaluating and improving agents on software engineering tasks.
@@ 2 Related Work
Multimodal and Computer-Use Agents. Recent works have explored using multimodal LLM agents to operate user interfaces such as web browsers (Koh et al., 2024; Deng et al., 2023; Zheng et al., 2024) and operating systems (Xie et al., 2024; Bonatti et al., 2024). Agent designs in these settings fall into two categories: (a) agents with predefined action sets (e.g., new_tab, go_back, click [element id]) that receive auxiliary information such as HTML accessibility trees (Yang et al., 2023a) for visual grounding; (b) pure computer-use agents operating with primitive keyboard and mouse actions, relying solely on screenshots (Anthropic, 2024; OpenAI, 2025; Qin et al., 2025). PwP supports evaluating both agent designs. Further, existing benchmarks such as OSWorld (Xie et al., 2024), AndroidWorld (Rawles et al., 2025), and WindowsAgentArena (Bonatti et al., 2024) evaluate agents on simple tasks like document editing and calendar management, leaving unclear whether performance on these tasks translates to complex, specialized domains like software engineering. PwP-Bench fills this gap by providing the first benchmark specifically designed to test whether computer-use agents can handle software engineering tasks. While some prior works explores specialized domains such as game playing (Tan et al., 2024) and a concurrent work explores scientific software (Sun et al., 2025), PwP is the first environment and PwP-Bench the first benchmark systematically evaluating computer-use agents for software engineering, a domain that is particularly noteworthy due to the presence of strong specialized agent baselines.
Software Engineering Agents. Software engineering agents have primarily relied on specialized scaffolding tailored to specific tools, languages, or tasks (Jin et al., 2024; Yang et al., 2024b). For instance, Agentless (Xia et al., 2024) uses Python-specific parsers, SWE-agent employs task-specific modifications (Abramovich et al., 2024; Yang et al., 2024b), and others depend on hand-engineered components like IPython kernels (Wang et al., 2024a) or custom browser views (Yang et al., 2024b). Our work takes a fundamentally different approach by evaluating whether computer-use agents –which interact through the same visual interface as human developers– can match these specialized agents. This also tests whether visual interaction with standard developer tools is sufficient for software engineering or if specialized APIs remain necessary. As PwP supports evaluating both designs, it enables direct comparison between computer-use and specialized agents across the diverse tasks in PwP-Bench, establishing a unified platform for understanding the capabilities and limitations of different agent designs. We refer readers to Appendix C for a more detailed related work.